# LRQ-DiT: Log-Rotation Post-Training Quantization of Diffusion Transformers for Text-to-Image Generation

### v1

@@ Abstract
Diffusion Transformers (DiTs) have achieved impressive performance in text-to-image generation. However, their high computational cost and large parameter sizes pose significant challenges for usage in resource-constrained scenarios. Post-training quantization (PTQ) is a promising solution to reduce memory usage and accelerate inference, but existing PTQ methods suffer from severe performance degradation under extreme low-bit settings. We identify two key obstacles to low-bit post-training quantization for DiT models: (1) model weights follow a Gaussian-like distribution with long tails, causing uniform quantization to poorly allocate intervals and leading to significant errors; (2) two types of activation outliers: (i) Mild Outliers with slightly elevated values, and (ii) Salient Outliers with large magnitudes concentrated in specific channels, which disrupt activation quantization. To address these issues, we propose LRQ-DiT, an efficient and accurate PTQ framework. We introduce Twin-Log Quantization (TLQ), a log-based method that aligns well with the weight distribution and reduces quantization errors. We also propose an Adaptive Rotation Scheme (ARS) that dynamically applies Hadamard or outlier-aware rotations based on activation fluctuation, effectively mitigating the impact of both types of outliers. We evaluate LRQ-DiT on PixArt and FLUX under various bit-width settings, and validate the performance on COCO, MJHQ, and sDCI datasets. LRQ-DiT achieves low-bit quantization of DiT models while preserving image quality, outperforming existing PTQ baselines.
@@ 1 Introduction
Figure 1: The overall framework of LRQ-DiT. (a) LRQ-DiT proposes the twin-log quantization (TLQ) method to address imbalanced weight distribution and excessive quantization errors. Compared to uniform quantization, TLQ significantly enhances the practicality of low-bit quantization. (b) LRQ-DiT utilizes the metric J to identify mild and salient outliers in activations, and proposes an adaptive rotation scheme that applies customized rotations to suppress the impact of each type of outlier.
@@
Diffusion models have achieved remarkable success in text-to-image generation, with recent advances expanding their applications across vision domains (Esser et al. 2024; Podell et al. 2023; Xie et al. 2024; Lin et al. 2025; Jiang et al. ; Zhou et al. 2024). Notably, DiTs replace the traditional U-Net backbone with the Transformer architecture, significantly improving scalability and computational efficiency. DiTs are becoming dominant in tasks such as image generation (Hatamizadeh et al. 2024; Chen et al. 2024b), video generation (Zhang et al. 2025), and image super-resolution. Despite their performance, the large number of parameters in DiT models poses serious challenges for deployment on resource-constrained edge devices, especially in terms of memory consumption and inference latency.
@@
Model compression (Liu et al. 2018; Hinton, Vinyals, and Dean 2015; Lin et al. 2024a) offers a promising acceleration solution, with quantization being one of the most effective techniques. By converting high-precision floating-point values into low-bit integers, quantization reduces both storage and computation costs, and has been widely adopted in large-scale Transformers (Yuan et al. 2022; Li et al. 2023b). PTQ has become increasingly popular for large models, such as LLMs and diffusion transformers, due to its training-free nature (Yao et al. 2022; Shao et al. 2023). Existing PTQ methods for DiTs typically focus on moderate configurations such as W4A8 and W8A8, but performance under lower-bit settings remains limited. PTQ4DiT (Wu et al. 2024a) and Q-DiT (Chen et al. 2025) incorporate time-step-specific activation patterns to achieve W4A8 quantization on DiT-XL/2, yet they exhibit performance degradation when applied to more advanced DiT variants such as PixArt and FLUX. ViDiT-Q (Zhao et al. 2024) proposes a dynamic quantization strategy, and SVDQuant (Li et al. 2024b) addresses outliers via kernel-level protection, both achieving strong performance at the 4 bit weight quantization. However, achieving accurate quantization for DiTs remains a significant challenge, particularly due to the difficulty in preserving precision for both weights and activations under extreme low-bit constraints.
@@
Our preliminary experiments identify two key challenges for low-bit quantization of DiT models: (1) Uniform quantizers fundamentally mismatch the Gaussian-like weight distribution of DiTs. We observe that almost all weight matrices in DiT models follow a Gaussian-like distribution with long tails on both sides. The Uniform quantization fails to allocate sufficient quantization intervals to the dense central region and ignores the long-tail extremes, resulting in large quantization errors under low-bit settings. (2) DiT activations contain two distinct types of outliers that disrupt low-bit quantization. The majority of outliers in activations are Mild Outliers, with values slightly exceeding normal activation levels and rarely surpassing 5. A smaller portion is Salient Outliers, represented as high-magnitude activations concentrated in specific channels. These outliers pose significant challenges for low-bit activation quantization.
@@
To address the challenges identified above, we propose LRQ-DiT, which integrates two core components: a twin-log quantization (TLQ) method for Gaussian-like model weights, and an adaptive rotation scheme (ARS) for accurate activations quantization. First, to better match the weight distribution, we introduce twin-log quantization. We apply a logarithmic transformation to model weights and quantize the positive and negative parts separately. A simple search-based clipping strategy is further employed to suppress extreme values in the long-tail regions. Second, we propose an adaptive rotation scheme to handle Mild and Salient activation outliers. To balance the accuracy and efficiency, we define an adaptive metric to measure the fluctuation level of activations across layers. Layers with low fluctuation are processed using Hadamard rotation, while layers with high fluctuation adopt dual transformations, combining greedy outlier-aware rotation with channel-wise permutation for enhanced precision. This design enables LRQ-DiT to effectively mitigate both types of outliers, while maintaining high precision and deployment efficiency.
Our main contributions are summarized as follows:
@@
  • Extensive quantization experiments conducted on PixArt and FLUX models under varying bit settings and datasets demonstrate the effectiveness of our method LRQ-DiT.
@@ 2.1 Text-to-Image Models
Text-to-image generation has made remarkable progress, largely driven by advancements in model architecture. Stable Diffusion (Rombach et al. 2022) bridges the semantic gap between text and image modalities, and its successor, Stable Diffusion 3 (Esser et al. 2024), further enhances multi-modal fusion. Subsequent models such as SDXL (Podell et al. 2023) and its variants provide a strong foundation for efficient and high-quality text-to-image generation. More recently, transformer-based architectures have replaced U-Net in diffusion models, opening new possibilities for downstream applications. PixArt-$\alpha$ (Chen et al. 2023) reduces training cost while maintaining image quality, and PixArt-$\Sigma$ (Chen et al. 2024a) significantly improves 4K image generation efficiency. FLUX (Labs 2024) (primarily FLUX.1-schnell and FLUX.1-dev) further demonstrates the capability of DiT-based models to generate high-resolution, high-fidelity images. In this work, we focus on diffusion transformers and propose LRQ-DiT, a low-bit weight–activation quantization framework tailored for DiT models to enhance the quality of text-to-image generation.
@@ 2.2 Model Quantization
Quantization reduces model memory by converting high-precision floating-point weights to low-bit integers, while activation quantization accelerates inference. Post-training quantization (PTQ) (Yang et al. 2024b, a) is also widely used in large-scale pre-trained models as it avoids costly retraining. For large language models (LLMs), outliers (Dettmers et al. 2022; Sun et al. 2024), the values significantly larger than most activations, pose a major challenge. SmoothQuant (Xiao et al. 2023) shifts quantization difficulty to weights for better 8-bit performance, while rotation-based methods (Ashkboos et al. 2024; Liu et al. 2024b; Lin et al. 2024b) balance outliers via rotation transformations, enabling competitive 4-bit weight–activation quantization. Applying these techniques to diffusion models is nontrivial due to time-step-dependent features. For diffusion models (DMs), Q-Diffusion (Li et al. 2023a) and PTQ4DM (Shang et al. 2023) considered the influence of time-steps when determining quantization parameters, achieving effective 8-bit quantization, which also stimulated subsequent works (He et al. 2023; Huang et al. 2024; Wang et al. 2024). However, there are still differences in the structure between DiTs and DMs. Recent works address these variations by assigning channel-wise quantization parameters (Chen et al. 2025), reallocating salient channels across time steps (Wu et al. 2024a), protecting activation outliers with low-rank branches (Li et al. 2024b), or using fine-grained grouping with dynamic quantization (Zhao et al. 2024). Preliminary progress has also been made in quantizing video generation models (Feng et al. 2025; Li et al. 2025). However, these methods still degrade significantly under extreme low-bit settings, motivating us to explore the underlying challenges and propose LRQ-DiT.
3 Method
3.1 Preliminaries
Quantization.
@@ 3.1 Preliminaries
Rotation Transformation.
Recently, some rotation-based methods (Ashkboos et al. 2024; Liu et al. 2024b; Lin et al. 2024b) eliminate outliers in LLM activations via the rotation matrix $\mathbf{R}$ along with corresponding weight transformations.
$\mathbf{Y}$ $=\mathbf{X}\mathbf{W}^{\top}=(\mathbf{X}\mathbf{R})(\mathbf{R}^{\top}\mathbf{W}^{\top}),$ (2)
where $~\mathbf{R}\mathbf{R}^{\top}=I\quad\text{and}\quad|\mathbf{R}|=1.$
@@ 3.2 Twin-Log Quantization for Weights
Motivation.
Inspired by previous studies (Li et al. 2025; Feng et al. 2025) and our experiments, we visualize the weight distributions of DiTs and observe that most weight matrices follow an approximately zero-mean Gaussian-like distribution with long tails on both sides, as shown in Figure 1(a). Existing works (Zhao et al. 2024; Wu et al. 2024a) primarily adopt uniform quantizers for weight quantization. However, our preliminary experiments reveal that 3-bit uniform quantization introduces a large error ($W_{q}-W||_{2}$) on PixArt weights, illustrated in Figure 2. This explains the severe accuracy degradation under low-bit settings. We attribute this to two factors: ➊ uniform quantizer cannot allocate sufficient intervals to the dense central region, limiting the expressive power of the majority of parameters; and ➋ uniform quantizer fails to handle the long-tail regions, where parameters are few but have large magnitudes.
@@
. To address the limitations of uniform quantization, we propose a twin-log quantization (TLQ) method designed to better fit the Gaussian-like weight distribution. Regarding issue ➊, we introduce a log-based quantizer to allocate more quantization intervals to the dense central region. Specifically, we first apply a logarithmic transformation to obtain $\bm{W}^{\prime}=\log_{2}(|\bm{W}|)$. The transformed weight matrix $\bm{W}^{\prime}$ is then divided into positive and negative parts using position masks $\bm{M^{+}}$ and $\bm{M^{-}}$:
@@
Search-based Clipping.
Inspired by learnable weight clipping (LWC) (Shao et al. 2023), we adopt a search-based clipping strategy to further suppress quantization errors in long-tail regions. Specifically, two hyperparameters, $\alpha$ and $\beta$, are introduced to clip the extreme values on both sides of the weight distribution, and their optimal values are determined via a simple grid search. The clipping-based scaling factors and zero-points are calculated as:
@@
We then introduce an integerization factor $2^{-I}$ (e.g., 1/64, 1/128, 1/256) (Wu et al. 2024b) to approximate the residual exponents by integers while minimizing quantization error:
@@
$\mathbb{I}^{X}$ and $s^{X}$ represent the integers and scales of activations. This formulation enables efficient hardware execution: the SHIFT operation accelerates power-of-two multiplications, and mask multiplications are implemented via fast bitwise AND operations. Therefore, TLQ can be deployed with lower latency and reduced computational cost.
@@ 3.3 Adaptive Rotation Scheme for Activations
Motivation.
We observe two distinct types of activation outliers in DiT models. Here, outliers refer to values significantly larger than typical activations. Based on their magnitude, we categorize them as  ➊ Mild Outliers and  ➋ Salient Outliers. As shown in Figure 1(b), most outliers slightly exceed typical activation values and rarely surpass 5, which we term Mild Outliers. In contrast, Salient Outliers appear in specific channels with much larger magnitudes. For instance, in the blocks.26.attn2.to_\_q layer of PixArt, there is a notable phenomenon of numerical outliers: 3.6% of the activation values exceed 10, while a small but critical subset (0.1%) surpasses 100, with the peak value reaching as high as 245. These disproportionately large activations pose significant challenges for low-bit weight-activation quantization. Besides, these Salient Outliers are prompt-independent and consistently observed across models such as PixArt and FLUX, with more pronounced patterns in PixArt. Notably, our observation differs from the Massive Outliers reported in LLMs (Sun et al. 2024; Liu et al. 2024a), which typically appear in a few specific tokens. In contrast, both Mild and Salient Outliers in DiT models are distributed across all tokens but confined to specific channels.
Model Precision Method COCO dataset MJHQ dataset sDCI dataset
@@
Table 1: Performance of LRQ-DiT (PixArt-$\Sigma$) text-to-image generation on different datasets. FID is calculated based on the images generated by FP16 inference. To ensure fairness in the comparison, we align the experimental setup and rerun SmoothQuant, QuaRot, ViDiT-Q, denoted as Method*. We follow the setting of ViDiT-Q to preserve FP precision for certain sensitive and parameter small layers (e.g., embed, norm_\_out, proj_\_out, adaln_\_single, caption_\_projection), denoted as [WnAn].
Adaptive Rotation Scheme (ARS)
. Recently, rotation-based transformations have been proven effective in mitigating outliers in LLMs. The use of Hadamard matrices $\mathbf{R}_{H}\in\{+1,-1\}^{n\times n}$ provides a simple and efficient way to smooth outliers (Ashkboos et al. 2024; Chee et al. 2023; Tseng et al. 2024). Moreover, DuQuant (Lin et al. 2024b) further enhances performance on massive outliers by constructing rotation matrices based on the largest outlier channel and introducing permutation transformations to balance the activation distribution. The Hadamard rotation is lightweight and straightforward, while DuQuant’s operations are more effective for handling high-magnitude outliers.
$\mathbf{J}$ $=\frac{\|\mathbf{X}\|_{F}}{\sqrt{BNC}}$ (10)
[…]
Model Precision Method COCO dataset MJHQ dataset sDCI dataset
@@
Ours  37.06  0.82 0.62 15.35  33.68  1.32 0.65 16.56  39.75  0.78 0.57 15.59
Table 2: Performance of LRQ-DiT (PixArt-$\alpha$) text-to-image generation on different datasets.
where $\mathbf{E}_{c^{*}}$ is a switch matrix exchanging the first and $c^{*}$ columns, and $\tilde{\mathbf{R}}$ is an initialized orthogonal matrix. We perform $n$ greedy updates on the top-$k$ influential channels to obtain $\hat{\mathbf{R}}_{1}$, which mitigates outliers more effectively than a fixed Hadamard matrix. To further balance inter-block outlier distributions, channels are reordered in a staggered manner by sorting their numerical ranges and alternately assigning them across blocks, denoted as an orthogonal matrix $\mathbf{P}$. After permutation, another rotation $\hat{\mathbf{R}}_{2}$ is applied to further smooth residual outliers, effectively improving low-bit quantization performance. Our adaptive rotation scheme effectively mitigates both types of outliers while balancing computational cost and quantization performance. Notably, the entire strategy is training-free, avoiding expensive optimization for hundreds of linear layers in DiT models. Using the calibration data, we apply additional rotations to about 5% of layers that exhibit Salient Outliers. Simply setting the threshold to 1 yields satisfactory results in practice.

### v2

# LRQ-DiT: Log-Rotation Post-Training Quantization of Diffusion Transformers for Image and Video Generation

@@ Abstract
Diffusion Transformers (DiTs) have achieved impressive performance in text-to-image and text-to-video generation. However, their high computational cost and large parameter sizes pose significant challenges for usage in resource-constrained scenarios. Effective compression of models has become a crucial issue that urgently needs to be addressed. Post-training quantization (PTQ) is a promising solution to reduce memory usage and accelerate inference, but existing PTQ methods suffer from severe performance degradation under extreme low-bit settings. After experiments and analysis, we identify two key obstacles to low-bit PTQ for DiTs: (1) the weights of DiT models follow a Gaussian-like distribution with long tails, causing uniform quantization to poorly allocate intervals and leading to significant quantization errors. This issue has been observed in the linear layer weights of different DiT models, which deeply limits the performance. (2) two types of activation outliers in DiT models: (i) Mild Outliers with slightly elevated values, and (ii) Salient Outliers with large magnitudes concentrated in specific channels, which disrupt activation quantization. To address these issues, we propose LRQ-DiT, an efficient and accurate post-training quantization framework for image and video generation. First, we introduce Twin-Log Quantization (TLQ), a log-based method that allocates more quantization intervals to the intermediate dense regions, effectively achieving alignment with the weight distribution and reducing quantization errors. Second, we propose an Adaptive Rotation Scheme (ARS) that dynamically applies Hadamard or outlier-aware rotations based on activation fluctuation, effectively mitigating the impact of both types of outliers. Extensive experiments on various text-to-image and text-to-video DiT models demonstrate that LRQ-DiT preserves high generation quality under low-bit weight–activation quantization, outperforming existing PTQ baselines.
@@ 1 Introduction
Figure 1: The overall framework of LRQ-DiT. (a) LRQ-DiT proposes the twin-log quantization (TLQ) method to address imbalanced weight distribution and excessive quantization errors. Compared to uniform quantization, TLQ significantly enhances the practicality of low-bit quantization. (b) LRQ-DiT utilizes the metric J to identify mild and salient outliers in activations, and proposes an adaptive rotation scheme (ARS) that applies customized rotations to suppress the impact of each type of outlier in activations.
@@
Diffusion models have achieved remarkable success in text-to-image and text-to-video generation tasks, with recent advances expanding their applications across vision domains [9, 33, 49, 56, 29, 26, 15, 65]. Notably, diffusion transformers (DiTs) replace the traditional U-Net backbone with the Transformer architecture, significantly improving scalability and computational efficiency. DiTs are becoming dominant in tasks such as image generation [3], video generation [59], and image super-resolution [17, 58]. Despite their performance, the large number of parameters in DiTs poses serious challenges for deployment on resource-constrained edge devices, especially in terms of memory consumption and inference latency.
@@
Model compression [51, 61, 37, 60, 24] offers a promising acceleration solution, with quantization being one of the most effective techniques. By converting high-precision floating-point values into low-bit integers, quantization reduces both storage and computation costs, and has been widely adopted in large-scale Transformers. Compared with quantization-aware training (QAT) [11, 47], PTQ has become increasingly popular for large models, such as Large Language Models (LLMs) [55, 36] and Diffusion Transformers (DiTs), due to its training-free nature. Existing PTQ methods for DiTs typically focus on moderate configurations such as W4A8 and W8A8, but performance under lower-bit settings remains limited. PTQ4DiT [43] and Q-DiT [6] incorporate time-step-specific activation patterns to achieve W4A8 quantization on DiT-XL/2, yet they exhibit performance degradation when applied to more advanced DiT variants such as PixArt and FLUX. ViDiT-Q [62] proposes a dynamic quantization strategy, and SVDQuant [20] addresses outliers via kernel-level protection, both improving the performance of W4A8 quantized models. Besides, some studies [10, 23] have made preliminary attempts on video generation models. However, achieving accurate quantization for DiTs remains a significant challenge, particularly due to the difficulty in preserving precision for both weights and activations under extreme low-bit constraints.
@@
Our preliminary experiments identify two key challenges for low-bit quantization of DiT models: (1) Uniform quantizers fundamentally mismatch the Gaussian-like weight distribution of DiTs. We observe that almost all weight matrices in DiT models follow a Gaussian-like distribution with long tails on both sides. The Uniform quantization fails to allocate sufficient quantization intervals to the dense central region and ignores the long-tail extremes, resulting in large quantization errors under low-bit settings. (2) DiT activations contain two distinct types of outliers that disrupt low-bit quantization. We discover that DiT models exhibit two distinct types of outliers. The majority of outliers in activations are Mild Outliers, with values slightly exceeding normal activation levels and rarely surpassing 5. A smaller portion is Salient Outliers, represented as high-magnitude activations concentrated in specific channels with their values approaching 300. These outliers pose significant challenges for low-bit activation quantization.
@@
To address the challenges identified above, we propose an accurate and efficient method LRQ-DiT, which integrates two core components: a twin-log quantization (TLQ) method for Gaussian-like model weights, and an adaptive rotation scheme (ARS) for accurate activations quantization. First, to better match the Gaussian distribution of weights, we introduce twin-log quantization. We apply logarithmic transformation to the model weights and quantize the positive and negative parts separately, allocating more quantization intervals to the dense areas in the middle. A simple search-based clipping strategy is further employed to suppress extreme values in the long-tail regions. Second, we propose an adaptive rotation scheme to handle Mild and Salient activation outliers. To balance the accuracy and efficiency, we define an adaptive metric to measure the fluctuation level of activations across layers. Layers with low fluctuation are processed using Hadamard rotation, while layers with high fluctuation adopt dual transformations, combining greedy outlier-aware rotation with channel-wise permutation for enhanced precision. This design enables LRQ-DiT to effectively mitigate both types of outliers, while maintaining high precision and deployment efficiency. Finally, comprehensive evaluations on PixArt and FLUX models for text-to-image generation, as well as OpenSORA model for text-to-video generation, prove the superiority of LRQ-DiT in achieving low-bit quantization without sacrificing generation quality. Our main contributions are summarized as follows:
@@
  • Extensive quantization experiments conducted on PixArt, FLUX and OpenSORA models under varying bit settings and different prompt sets demonstrate the effectiveness and robustness of LRQ-DiT, especially in low-bit scenarios.
@@ 2.1 Text-to-Image Models
Text-to-image generation has made remarkable progress, largely driven by advancements in model architecture. Stable Diffusion [34] bridges the semantic gap between text and image modalities, and its successor, Stable Diffusion 3 [9], further enhances multi-modal fusion. Subsequent models such as SDXL [33] and its variants provide a strong foundation for efficient and high-quality text-to-image generation. More recently, transformer-based architectures have replaced U-Net in diffusion models, opening new possibilities for downstream applications. PixArt-$\alpha$ [4] reduces training cost while maintaining image quality, and PixArt-$\Sigma$ [5] significantly improves 4K image generation efficiency. FLUX [18] (primarily FLUX.1-schnell and FLUX.1-dev) further demonstrates the capability of DiT-based models to generate high-resolution, high-fidelity images. OpenSORA [63] brings DiTs to the video domain with a real-time capable, multimodal generation pipeline. However, the parameter and computational complexity of the model are constantly increasing, which poses difficulties for further development and deployment.
2.2 Text-to-Video Models
Given that DiT has achieved remarkable success and garnered significant attention in text-to-image generation tasks, numerous studies have attempted to introduce it into text-to-video generation tasks to further expand its range of applications. The Latte [32] model achieves efficient and high-quality video generation through its latent diffusion transformer architecture. Sora [30] has been trained for generating clear and realistic scene videos from textual instructions. OpenSORA [63] employs the Spatial-Temporal Diffusion Transformer (STDiT) technology to accomplish a variety of visual generation tasks, facilitating the synthesis of high-definition videos. CogVideoX [54] employs progressive training with multi-resolution frame packing to generate high-quality, coherent, and long-duration videos. HunyuanVideo [16] trains a video generation model with over 13 billion parameters. Wan [41] has made substantial progress in generation by introducing innovative methods based on the DiT framework. These works have greatly expanded the applications of DiT models and achieved outstanding results. However, as the number of model parameters and computational complexity continue to increase, it also presents significant challenges for further applications. This issue is particularly evident in the text-to-video generation task.
@@ 2.3 Model Quantization
[…]
3 Preliminaries
3.1 Quantization
@@
3.2 Rotation Transformation
Recently, some rotation-based methods [1, 31, 25] eliminate outliers in LLM activations via the rotation matrix $\mathbf{H}$ along with corresponding weight transformations. This transformation improves the quantization performance of the model while ensuring computational invariance.
$\mathbf{Y}$ $=\mathbf{X}\mathbf{W}^{\top}=(\mathbf{X}\mathbf{H})(\mathbf{W}\mathbf{H})^{\top},$ (2)
where $~\mathbf{H}\mathbf{H}^{\top}=I\quad\text{and}\quad|\mathbf{H}|=1.$
$\mathbf{H}$ is typically a Hadamard matrix, which is orthogonal and has elements taking values in $\{+1,-1\}$. A Hadamard matrix is a square matrix of size $2^{n}$, as shown in the Equation (3).
$\mathbf{H}_{2}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&1\\ 1&-1\end{bmatrix}\quad\text{and}\quad\mathbf{H}_{2^{n}}=\mathbf{H}_{2}\otimes\mathbf{H}_{2^{n-1}}.$ (3)
4 Method
@@ 4.1 Twin-Log Quantization for Weights
Motivation
Inspired by previous studies [23, 10] and our experiments, we visualize the weight distributions of DiTs and observe that most weight matrices follow an approximately zero-mean Gaussian-like distribution with long tails on both sides, as shown in Figure 1(a). Existing works [62, 43] primarily adopt uniform quantizers for weight quantization. However, our preliminary experiments reveal that 3-bit uniform quantization introduces a large error ($W_{q}-W||_{2}$) on PixArt weights, illustrated in Figure 2. This explains the severe accuracy degradation under low-bit settings. We attribute this to two factors: ➊ uniform quantizer cannot allocate sufficient intervals to the dense central region, limiting the expressive power of the majority of parameters; and ➋ uniform quantizer fails to handle the long-tail regions, where parameters are few but have large magnitudes.
@@
To address the limitations of uniform quantization, we propose a twin-log quantization (TLQ) method designed to better fit the Gaussian-like weight distribution. Regarding issue ➊, we introduce a log-based quantizer to allocate more quantization intervals to the dense central region. Specifically, we first apply a logarithmic transformation to obtain $\bm{W}^{\prime}=\log_{2}(|\bm{W}|)$. The transformed weight matrix $\bm{W}^{\prime}$ is then divided into positive and negative parts using position masks $\bm{M^{+}}$ and $\bm{M^{-}}$:
@@
Search-based Clipping
Inspired by learnable weight clipping (LWC) [36], we adopt a search-based clipping strategy to further suppress quantization errors in long-tail regions. Specifically, two hyperparameters, $\alpha$ and $\beta$, are introduced to clip the extreme values on both sides of the weight distribution, and their optimal values are determined via a simple grid search. The clipping-based scaling factors and zero-points are calculated as:
@@
We then introduce an integerization factor $2^{-I}$ (e.g., 1/64, 1/128, 1/256) [44] to approximate the residual exponents by integers while minimizing quantization error:
@@
$\mathbb{I}^{X}$ and $s^{X}$ represent the integers and scales of activations. This formulation enables efficient hardware execution: the SHIFT operation accelerates power-of-two multiplications, and mask multiplications are implemented via fast bitwise AND operations. Therefore, TLQ can be deployed with lower latency and reduced computational cost, while maintaining high quantization accuracy. We provide the speedup measurement in Section 5.5.
@@ 4.2 Adaptive Rotation Scheme for Activations
Motivation
We observe two distinct types of activation outliers in DiT models. Here, outliers refer to values significantly larger than typical activations. Based on their magnitude, we categorize them as  ➊ Mild Outliers and  ➋ Salient Outliers. As shown in Figure 1(b), most outliers slightly exceed typical activation values and rarely surpass 5, which we term Mild Outliers. In contrast, Salient Outliers appear in specific channels with much larger magnitudes. For instance, in the blocks.26.attn2.to_\_q layer of PixArt, there is a notable phenomenon of numerical outliers: 3.6% of the activation values exceed 10, while a small but critical subset (0.1%) surpasses 100, with the peak value reaching as high as 245. These disproportionately large activations pose significant challenges for low-bit weight-activation quantization. Besides, these Salient Outliers are prompt-independent and consistently observed across models such as PixArt and FLUX, with more pronounced patterns in PixArt.
Notably, our observation differs from the Massive Outliers reported in LLMs [38, 28], which typically appear in a few specific tokens. In contrast, both Mild and Salient Outliers in DiT models are distributed across all tokens but confined to specific channels.
Adaptive Rotation Scheme (ARS)
Recently, rotation-based transformations have been proven effective in mitigating outliers in LLMs. The use of Hadamard matrices $\mathbf{R}_{H}\in\{+1,-1\}^{n\times n}$ provides a simple and efficient way to smooth outliers [1, 2, 39]. Moreover, DuQuant [25] further enhances performance on massive outliers by constructing rotation matrices based on the largest outlier channel and introducing permutation transformations to balance the activation distribution. The Hadamard rotation is lightweight and straightforward, while DuQuant’s operations are more effective for handling high-magnitude outliers.
$\mathbf{J}$ $=\frac{\|\mathbf{X}\|_{F}}{\sqrt{BNC}}$ (11)
[…]
Table 1: Performance of LRQ-DiT (PixArt-$\Sigma$) text-to-image generation on different prompt sets. FID is calculated based on the images generated by FP16 inference. To ensure fairness in the comparison, we align the experimental setup and rerun SmoothQuant, QuaRot, DuQuant, ViDiT-Q, denoted as Method*. We follow the setting of ViDiT-Q to preserve FP precision for certain sensitive and parameter small layers (e.g., embed, norm_\_out, proj_\_out, adaln_\_single, caption_\_projection), denoted as [WnAn]. These settings also apply to other Tables.
Model Precision Method COCO prompt set MJHQ prompt set sDCI prompt set
@@
where $\mathbf{E}_{c^{*}}$ is a switch matrix exchanging the first and $c^{*}$ columns, and $\tilde{\mathbf{R}}$ is an initialized orthogonal matrix. We perform $n$ greedy updates on the top-$k$ influential channels to obtain $\hat{\mathbf{R}}_{1}$, which mitigates outliers more effectively than a fixed Hadamard matrix. To further balance inter-block outlier distributions, channels are reordered in a staggered manner by sorting their numerical ranges and alternately assigning them across blocks, denoted as an orthogonal matrix $\mathbf{P}$. After permutation, another rotation $\hat{\mathbf{R}}_{2}$ is applied to further smooth residual outliers, effectively improving low-bit quantization performance. Our adaptive rotation scheme effectively mitigates both types of outliers while balancing computational cost and quantization performance. The adaptive rotation scheme is particularly effective for DiT models, as it handles the two types of outliers in different ways to balance efficiency and performance. Notably, the entire strategy is training-free, avoiding expensive optimization for hundreds of linear layers in DiT models. Using the calibration data, we apply additional rotations to about 5% of layers that exhibit Salient Outliers. Simply setting the threshold to 1 yields satisfactory results in practice. The resulting computational overhead during inference is negligible compared to the performance benefits achieved. We can achieve satisfactory results by simply setting the threshold to 1. With the help of calibration data, we only perform additional processing on 5% of the layers that contain salient outliers. Compared to the benefits gained, the overhead introduced during the inference phase is negligible.
Table 2: Performance of LRQ-DiT (PixArt-$\alpha$) text-to-image generation on different prompt sets.
Model Precision Method COCO prompt set MJHQ prompt set sDCI prompt set
@@
Ours  37.06  0.83 0.62 15.35  33.68  1.32 0.65 16.56  39.75  0.78 0.57 15.59
4.3 Theoretical Analysis
Lemma 1.
Let $x\in\mathbb{R}^{C}$ be an activation vector and $R\in\mathbb{R}^{C\times C}$ be an orthogonal rotation matrix. The transformed vector $x^{\prime}=xR$ preserves the L2-norm:
$\|x^{\prime}\|_{2}=\|x\|_{2}.$
Furthermore, if $x$ is dominated by a salient outlier such that its L-infinity norm is $\|x\|_{\infty}$, the transformed vector’s L-infinity norm is suppressed as:
$\|x^{\prime}\|_{\infty}\approx\frac{\|x\|_{\infty}}{\sqrt{C}}$
Theorem 1.
Let $X$ be a C-dim activation tensor containing a salient outlier of magnitude $\alpha=\|X\|_{\infty}$. Let $\textrm{MSE}_{\textrm{before}}=\textrm{MSE}(\textrm{TLQ}(X))$ and $\textrm{MSE}_{\textrm{after}}=\textrm{MSE}(\textrm{TLQ}(\textrm{ARS}(X)))$ denote the mean squared error of Twin-Log Quantization before and after the application of ARS, respectively. Then, the ratio of these errors is given by:
$\frac{\textrm{MSE}_{\textrm{after}}}{\textrm{MSE}_{\textrm{before}}}\approx(1-\frac{log_{2}\sqrt{C}}{log_{2}\alpha})^{2}.$
Proof.
We first map the input $x$ into the logarithmic domain via $y=log_{2}(|x|)$, where uniform quantization with step size $\Delta_{y}$ introduces an error $e_{y}$. Mapping back to the linear domain, the quantized value is $x_{q}=|x|\cdot 2^{e_{y}}$, leading to a linear-domain error,
$e_{x}=x_{q}-|x|=|x|(2^{e_{y}}-1)\approx|x|(ln2)e_{y},$
where the approximation follows from the first-order expansion of $2^{e_{y}}$. Hence the mean squared error is,
$\textrm{MSE}_{x}=(|x|ln2)^{2}\sigma_{e_{y}}^{2},$
showing the quantization error scales quadratically with the input amplitude and grows rapidly in the presence of outliers.
[…]
Therefore, the total MSE before and after ARS are,
$\textrm{MSE}_{\textrm{before}}\propto(log_{2}\alpha)^{2}\|X\|_{F}^{2},\textrm{MSE}_{\textrm{after}}\propto(log_{2}\frac{\alpha}{\sqrt{C}})^{2}\|X\|_{F}^{2},$
and their ratio simplifies to,
$\frac{\textrm{MSE}_{\textrm{after}}}{\textrm{MSE}_{\textrm{before}}}\approx(1-\frac{log_{2}\sqrt{C}}{log_{2}\alpha})^{2},$
which quantifies the error reduction due to the shrinkage of the effective dynamic range by ARS. ∎
This result quantitatively demonstrates the powerful synergy between ARS and TLQ. It reveals that the error reduction becomes more pronounced as the vector dimensionality, C, increases. ARS leverages high dimensionality to effectively eliminate large outliers, reshaping the data into a more compact, low-magnitude region. By removing the sparse, high-magnitude tail, ARS allows TLQ’s precision-allocation mechanism to operate with maximum efficiency, leading to a substantial and scalable reduction in quantization error.